### v4

# AAppendix References

Abstract
A key strategy in societal adaptation to climate change is using alert systems to prompt preventative action and reduce the adverse health impacts of extreme heat events. This paper implements and evaluates reinforcement learning (RL) as a tool to optimize the effectiveness of such systems. Our contributions are threefold. First, we introduce a new publicly available RL environment enabling the evaluation of the effectiveness of heat alert policies to reduce heat-related hospitalizations. The rewards model is trained from a comprehensive dataset of historical weather, Medicare health records, and socioeconomic/geographic features. We use scalable Bayesian techniques tailored to the low-signal effects and spatial heterogeneity present in the data. The transition model uses real historical weather patterns enriched by a data augmentation mechanism based on climate region similarity. Second, we use this environment to evaluate standard RL algorithms in the context of heat alert issuance. Our analysis shows that policy constraints are needed to improve RL’s initially poor performance. Third, a post-hoc contrastive analysis provides insight into scenarios where our modified heat alert-RL policies yield significant gains/losses over the current National Weather Service alert policy in the United States.
Codehttps://github.com/NSAPH-Projects/heat-alerts˙RL
Simulatorhttps://github.com/NSAPH-Projects/weather2alert
Extended version (with appendices)https://arxiv.org/abs/2312.14196
1 Introduction
Extensive evidence links exposure to extreme heat to increases in morbidity and mortality (Ebi et al. 2021). Heat alerts are a practical and low-cost intervention to mitigate these effects (Ebi et al. 2004) by encouraging protective measures such as hydrating more, avoiding physical exertion outdoors, and opening cooling centers. However, studies investigating the effectiveness of heat alerts have observed mixed results (Weinberger et al. 2018; Wu et al. 2023). Developing methods to optimize the issuance of heat alerts for public health is an open problem.
Various challenges stand in the way of solving this problem. First, issuing too many heat alerts may lead to alert fatigue (Nahum-Shani et al. 2017) and health-protective resource depletion on both individual and community/institutional levels. Sequential decision-making (SDM) methods offer a promising yet unexplored avenue for tackling this issue. Second, despite representing a significant health threat on the population scale, local health impacts of heat—and therefore heat alerts—are small and easily confounded in observational data sets (Weinberger et al. 2021). Rare events and low signal (small effects) have been shown to challenge algorithmic decision-making (Frank, Mannor, and Precup 2008; Romoff et al. 2018). Third, mainstream SDM methods are not suitable for spatially heterogeneous settings in which a single policy is not equally effective in all regions or dynamically changing contexts (Padakandla 2021). However, attempting to identify independent policies for each location drastically reduces the amount of available historical data.
In this paper, we lay the foundation for addressing these challenges by introducing a framework for optimizing heat alert issuance using reinforcement learning (RL). RL allows learning SDM policies for determining when to issue heat alerts, aiming to minimize the population health risk as a negative reward signal. The ultimate vision for this vein of research is deploying data-driven policies to alert the public about extreme heat events (and, eventually, other environmental exposures such as extreme cold or wildfire smoke). In addition to our analysis breaking ground towards this goal, a variety of follow-up work is facilitated by the publication of our SDM environment in an open-source Python package, weather2alert [^1] https://github.com/NSAPH-Projects/weather2alert, compatible with the Gymnasium framework for RL (Towers et al. 2023). This challenging, data-driven environment can be used as a benchmark. Further, our methodology to create this environment can be used as a blueprint for other RL simulators, particularly for problems with exogenous state space components and multiple locations or individuals.
Figure 1: Overview of the heat alerts RL framework.
The two main components of our heat alerts RL framework are visualized in Figure 1 and summarized below. First, we create a realistically challenging SDM environment for heat alert issuance, structured as a Markov Decision Process (MDP) under a budget constraint. Key components of the dataset used to create this environment are heat alerts issued by the U.S. National Weather Service (NWS), hospitalization records from Medicare, ambient heat index, and other covariates informed by the epidemiological literature. To estimate the rewards under both observed (NWS) and counterfactual alert policies, we fit a statistical model of hospitalizations that allows for spatiotemporal heterogeneity, uncertainty quantification, and fast inference using variational Bayes (Tzikas, Likas, and Galatsanos 2008). For the state-transition model, we exploit a factorization of the state space into an endogenous component with known transition dynamics and an exogenous component sourced from observed weather trajectories. This use of real data allows us to make grounded inferences about the effectiveness of both observed and counterfactual heat alert policies in the U.S. We call this environment structure Bayesian Rewards Over Actual Climate History (BROACH).
Second, we use our environment to train RL models and then to evaluate the RL policies as well as the NWS and alternative baseline policies—all subject to the same alert budget. Our evaluation identifies issues during training that lead to poor performance of standard, widely-used RL algorithms. We address these issues by introducing conceptually simple modifications which enable learning of policies that reduce hospitalizations relative to the NWS policy. These modifications include restricting the RL model to issue alerts only on extremely hot days and optimizing policies for each location separately. Lastly, we perform contrastive explanation (van der Waa et al. 2018; Narayanan, Lage, and Doshi-Velez 2022) of the RL policies by comparing their attributes (e.g., how often does the policy issue alerts on consecutive days) within and across locations to those of the NWS and alternative baseline policies. Specifically, we use visualizations and Classification and Regression Trees (CART) to illustrate systematic differences between these policies and identify characteristics of locations where RL-based policies might offer the greatest improvements in public health.
2 Related Work
Heat alert optimization
Recent approaches to improving the issuance of heat alerts include (i) the development of a causal inference technique for stochastic interventions to infer whether increasing the probability of issuing a heat alert would be beneficial (Wu et al. 2023) and (ii) a comparison of methods to identify optimal thresholds above which heat alerts should always be issued (Masselot et al. 2021). Crucially, neither of these approaches addresses the complications of sequential dependence, i.e., the potential for alert fatigue and running out of resources to deploy precautionary measures. To our knowledge, SDM techniques have not yet been explored in a climate and health setting.
RL with exogenous states
Several recent papers utilize a decomposition of the state space into exogenous vs. endogenous components, but none of these methods apply directly in our setting. Efroni et al. (2022) provide theoretical analysis and guarantees for sample efficiency in tabular (finite state) environments with an unknown subset of exogenous variables. Sinclair et al. (2023) introduce “hindsight learning” for situations with a known reward function and an exact solver under a fixed exogenous trajectory. Lee et al. (2023) also use hindsight, but do so in the context of identifying latent states within a partially observable MDP. Contrary to Levine, Stone, and Zhang (2024) and Efroni et al. (2024), we do not consider exogenous states as distractors since weather patterns are crucial for decision-making.
Statistical modeling for RL environments
Using a statistical model for the reward function in an RL environment is common practice in settings with empirical data. Bayesian models in particular have been used to identify the effect of intervention in health-related applications and dynamic treatment regimes (Liao et al. 2020; Tec, Duan, and Müller 2023; Zajonc 2012). For spatiotemporal or otherwise heterogeneous environments, previous studies have also used a combination of local modeling and global modeling that takes into account the spatial or hierarchical structure to create simulated environments for RL (Wu et al. 2021; Li, Zheng, and Yang 2018; Agarwal et al. 2021).
Constrained learning
The use of domain knowledge-based policy restrictions in RL has been found to speed up learning and to guide the learned policy in pragmatic directions (Mu et al. 2021). This supports our restriction of RL heat alerts to very hot days. More generally, incorporating “cost” or “safety” constraints is of interest for many RL applications (Laber et al. 2018; Carrara et al. 2019). A common approach in the constrained RL literature is to use Lagrange multipliers (Guin and Bhatnagar 2023; Ray, Achiam, and Amodei 2019), however, these methods do not enforce a strict constraint/budget. To enable direct comparison with the observed NWS policy, we impose a strict alert budget, which can be viewed as a simplified variation of the indicator approach formulated by Xu, Zhan, and Zhu (2022). An alternative approach to handling both the alert budget and restriction of alerts to very hot days would be to model the extent of distributional overlap with the observed NWS policy, and penalize actions that have too little overlap (Xu, Zhan, and Zhu 2022). However, our approach has the benefit of being relatively interpretable.
Contrastive policy explanations
There is a large literature on post-hoc explainability in RL (Heuillet, Couthouis, and Díaz-Rodríguez 2021). Contrastive analysis has been used to explain differences between RL policies and more familiar / intuitive policies (van der Waa et al. 2018; Narayanan, Lage, and Doshi-Velez 2022). We note that while CART has been used to try and mimic RL policies (Puiutta and Veith 2020), our use of it to analyze differences between policies has not been previously documented.
3 Problem Setup
3.1 RL Preliminaries
[…]
$J(\pi):=\mathbb{E}_{s_{0}\sim d_{0}}[V_{0}^{\pi}(s_{0})].$ (1)
Under regularity conditions (Puterman 2014), the optimal policy is deterministic and the unique solution (up to ties) of the Bellman optimality equation
$Q^{*}_{t}(s,a)=R(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim P(s,a)}[\max_{a^{\prime}}Q^{*}_{t+1}(s^{\prime},a^{\prime})],$ (2)
where $Q^{*}_{t}$ denotes the state-action value function of $\pi^{*}_{t}$. The optimal action satisfies $\pi^{*}_{t}(s)=\arg\max_{a}Q^{*}_{t}(s,a)$.
RL algorithms search for the optimal policy of an MDP by interacting with an environment that generates rewards and transitions, where the RL algorithm does not need knowledge of the transition function (Sutton and Barto 2018).
3.2 Issuing Heat Alerts as a Constrained MDP
The heat alerts MDP components are summarized here. See also Table S2 of Appendix A    Heat Alerts MDP Terminology. Each MDP ${\mathcal{M}}_{k}$ corresponds to a geographic area $k\in{\mathcal{K}}$, and episodes are indexed as $j\in{\mathcal{J}}$. In this work, geographic areas are U.S. counties, and each episode spans the warm season / summer from May 1st to September 30th of a specific year ($H=152$ days).
The action $a_{t}$ on day $t$ is either 1 (issue a heat alert) or 0 (do not issue a heat alert). To mitigate alert fatigue, we adopt a strict action budget-constrained approach, optimizing the policy subject to $\sum_{t=0}^{H-1}a_{t}\leq b$.
The reward $r_{t}=R(s_{t},a_{t})$ is the expected rate of heat-related hospitalizations at time $t$, transformed such that fewer hospitalizations correspond to a greater reward. Specifically, if $\rho_{t}(a_{t}):=\rho(s_{t},a_{t})$ represents the per capita rate of heat-related hospitalizations on day $t$ if action $a_{t}\in\{0,1\}$ is taken at state $s_{t}$, then $r_{t}\propto-\rho_{t}(a_{t})+\text{const}$.
The state vector $s_{t}$ contains the factors underlying the effect of heat on hospitalizations and the effectiveness of heat alerts at reducing hospitalizations. It can be decomposed as $s_{t}=(\xi_{t},x_{t})$, where $\xi_{t}$ is the exogenous component and $x_{t}$ is the endogenous component. The exogenous component is defined as the portion of the state space that is not influenced by the agent’s actions $a_{t}$, such as the weather. The transition function of the endogenous component (heat alert history) is known and deterministic. Hence, the full state transition function admits a factorization
$P(s_{t+1}|s_{t},a_{t})$ $=P((\xi_{t+1},x_{t+1})|(\xi_{t},x_{t}),a_{t})$ (3)
$=P_{\xi}(\xi_{t+1}|\xi_{t})P_{x}(x_{t+1}|x_{t},a_{t})$
4 BROACH: An RL Environment for Optimizing Heat Alert Issuance
In creating an interactive heat alert environment, our design objectives are (i) allow use of general-purpose RL algorithms, (ii) be grounded in real data, and (iii) handle the low signal and heterogeneity of health impacts of heat and heat alerts across space and time. To meet these objectives, we introduce a methodology termed Bayesian Rewards Over Actual Climate History (BROACH). BROACH can be generalized to other climate-related events and interventions.
4.1 Data Sources
Here we provide an overview of the data used in the RL environment; Appendix B    Details of the Data Set contains more details.
Heat alerts and heat index
We use the dataset of Weinberger et al. (2021) with daily, county-level records (2006-2016, May-September) of heat alerts issued by the NWS, as well as heat index, which is a measure of the combined effect of temperature and humidity. To understand the observed heat alert data, note that while the decision to issue an alert is based on temperature thresholds, it is also strongly affected by the discretion of the local office (Hawkins, Brown, and Ferrell 2017). Analysis by (Hondula et al. 2022) suggests that spatial variability in the current NWS/local office approach is not well aligned with the health risk from heat.
Heat-related hospitalizations
We merged the heat alert and weather data with daily, county-level hospitalization counts from Medicare. We included cause-specific hospitalizations previously found to be associated with extreme heat in the Medicare population: septicemia, peripheral vascular disease, urinary tract infections, and diabetes mellitus with complications (Bobb et al. 2014). We excluded heat stroke and fluid and electrolyte disorders (which were also found to be associated with extreme heat in Medicare) because Weinberger et al. (2021) observed a positive association between heat alerts and hospitalizations with these causes. They hypothesized it was due to increased awareness of heat-related symptoms and seeking medical care. Figure S4 illustrates this scenario as an unobserved mediation problem (Pearl 2012), which affects the identification of the preventive effect of heat alerts. We refer to the remaining causes, which we pool together into one outcome variable, as not-obviously heat-related (NOHR) hospitalizations.
Counties
We consider the 761 counties with a population greater than 65,000 to avoid locations with very few hospitalizations and to focus on the most populous areas. Figure 2 shows the counties considered, which account for approximately 75% of the population and 25% of the number of counties.
Figure 2: Map of the counties considered and their regional climate zone classifications. All the colored counties are used in the Bayesian rewards model and RL environment; the 30 counties with annotated FIPS codes are used in the RL experiments.
County-level characteristics
Spatial heterogeneity in the health effects of heat and heat alerts can occur due to people in different regions being affected differently by the same absolute temperatures due to geographical self-selection and climate adaptation, differential health susceptability and/or individual agency associated with socioeconomic status and population density, and variable response to heat alerts due to local policy and political ideology (Ng et al. 2014; Zanobetti et al. 2013; Errett et al. 2023; Cutler et al. 2018). To characterize this spatial heterogeneity, we compiled a set of county-level covariates: population density and median household income (U.S. Census Bureau 2014), regional classifications of climate zones (U.S. Energy Information Administration 2020), broadband (internet) usage data (Microsoft AI for Good Research Lab 2021), presidential election returns (MIT Election Data and Science Lab 2018), and fine particulate matter (PM2.5) (Hammer et al. 2020). For the latter, there is strong evidence of adverse synergistic health impacts of air pollution and heat (Anenberg et al. 2020).
4.2 The State Space
Recall that the state space can be factorized as $s_{t}=(\xi_{t},x_{t})$, where $\xi_{t}$ is the exogenous component and $x_{t}$ is the endogenous component. The exogenous component contains the quantile value of the day’s observed heat index (QHI)—within each county across warm seasons for all the years—which allows us to account for long-term climate adaptation. It also contains time-varying factors which can modify the heat-health relationship (Heo, Bell, and Lee 2019; Anderson and Bell 2011): day of summer (which is equivalent to $t$), weekend status, and excess heat compared to the last three days. The endogenous component contains the number of heat alerts issued in the last 14 days, an indicator of whether a heat alert was issued yesterday, and the remaining alert budget for that episode.
4.3 The Reward Function: A Hierarchical Model of Hospitalizations
To learn the expected hospitalization rate, and thereby the transformed reward $r_{t}$, we train a Bayesian hierarchical model with the desiderata of (a) facilitating obtainment of confidence intervals when evaluating RL policies and (b) allowing for spatial heterogeneity in the health effects of heat and heat alerts. Let $\rho_{t}^{(k,j)}(a)$ be the expected per capita NOHR hospitalization rate at time $t$ in county-summer $(k,j)$ when taking action $a$. We propose the scaled reward function $r_{t}^{(k,j)}(a):=C_{1}-C_{2}\rho_{t}^{(k,j)}(a)$ where $C_{1}$ and $C_{2}$ are positive scaling constants. (Note that choices of $C_{1}$ and $C_{2}$ do not change the optimal policy, but can help stabilize learning.) We choose $C_{1}=1$ and $C_{2}$ as the reciprocal of the observed hospitalization rate through the entire summer.
Let $y_{t}^{(k,j)}(a)$ denote the counterfactual outcome representing the number of NOHR hospitalizations for action $a\in\{0,1\}$, and $n^{(k,j)}$ denote the total population susceptible to hospitalization in county $k$ in episode $j$. We introduce two functions—to be learned—governing the counterfactual hospitalization rate in county $k$ as a function of $s_{t}^{(k,j)}$. First, $\lambda_{k}\colon S\to\mathbb{R}^{+}$ denotes the baseline rate of hospitalizations when no alert is issued. Second, $\tau_{k}\colon S\to(0,1)$ expresses the effectiveness of issuing an alert as a multiplicative reduction factor. The top-level hospitalization model is
$y_{t}^{(k,j)}(a)$ $\sim\text{Poisson}(n^{(k,j)}\rho_{t}^{(k,j)}(a)),$ (4)
$\rho_{t}^{(k,j)}(a)$ $:=\lambda_{k}(s_{t}^{(k,j)})(1-a\cdot\tau_{k}(s_{t}^{(k,j)})),$
for all $(k,j)\in{\mathcal{K}}\times{\mathcal{J}}$ and $t\in\{0,\ldots,H-1\}$. The Poisson distribution/loss function is the natural choice for count data and recommended for health outcomes data due to its correspondence to the Cox survival model (Austin 2017).
To estimate the functions $\lambda_{k}$ and $\tau_{k}$ under limited data and low health signal, we incorporate a data-driven random effects prior (using the county-level features) and inject domain knowledge on the signs of certain coefficients. We train this model using variational inference, using the Python package Pyro (Harpole et al. 2019). The details of this approach, as well as model diagnostics, are in Appendix C    Details of the Bayesian Rewards Model. Note that the fitted coefficients in the alert effectiveness component indicate the presence of alert fatigue. To utilize this rewards model within the BROACH environment, we draw from the variational posterior distribution at the start of each episode, and then use those coefficients to calculate $r^{(k,j)}_{t}$ given $\big{(}s^{(k,j)}_{t},a^{(k,j)}_{t}\big{)}$ for $t=0,...,H-1$.
4.4 The Alert Budget
Each episode in our MDP starts with a fixed alert budget $b$. This can be specified as a constant, sampled from a range of values, or set to the observed (NWS) value. As the agent steps through the episode, once $\sum_{h=0}^{t}a_{h}=b$, it is prohibited from issuing any more alerts: it can only obtain the reward $r_{h}^{(k,j)}(0)$ for days $h=t+1,...,H-1$.
4.5 The Transition Function: Observing Actual Weather/Climate History
We leverage the decomposition equation 3 to use real data for the exogenous trajectories, and introduce a data augmentation scheme in the context of single-county RL.
Sampling of exogenous trajectories
Recall that the endogenous aspect of our state space, heat alert history, is deterministically updating, so does not need to be modeled probabilistically. The key observation in the heat alert setting is that the more complicated aspect of the environment, the weather/climate, also does not need to be modeled because it is completely exogenous to heat alert decision-making. In this setting, developing a model for the weather is not only unnecessary, but would unavoidably introduce error. Instead, we sample real observed weather trajectories.
Regional data augmentation
The issue with this approach is that there are only 11 years of exogenous trajectories (2006-2016) available per county during which we also have heat alert data—which is needed when we use observed alert budgets (paired with observed weather) for comparison with the NWS. To mitigate the potential for RL overfitting to these 11 years, we propose data augmentation by sampling exogenous trajectories from other counties in the same regional climate zone [^4] Due to the large size/geographic range of the “Cold” zone, we grouped its counties into eastern and western subsets. (U.S. Energy Information Administration 2020) during RL training and validation / tuning.
5 Learning and Evaluation
We conduct experiments to (a) assess the ability of standard RL algorithms to learn effective heat alert policies in the BROACH environment, (b) test various modifications to these algorithms, and (c) compare the resulting policies to the NWS policy and other intuitive baselines. We then investigate heterogeneous performance of the RL policies to provide domain-relevant insights. Code for all steps is available in our GitHub repository [^5] https://github.com/NSAPH-Projects/heat-alerts_RL.
5.1 Policy Constraint: Very Hot Days
Foreshadowing our findings, standard RL algorithms struggle to learn to conserve their alert budgets for later in the summer, resulting in poor performance. A major modification we implement to encourage more effective behavior is restricting the issuance of heat alerts to days above a QHI threshold, optimized separately for each county. To optimize the QHI threshold, we test the sequence of values between 0.5 and 0.9 (ensuring overlap with the NWS policy), by every 0.05, and select the value that yields the best return on our validation set (specified in the following section). Names of models including this adaptation have the suffix “.QHI”.
5.2 Experimental Setup
For all training and evaluations, to directly compare counterfactual alert policies with the NWS policy, we fix $b$ in each county-summer to the observed number of heat alerts.
RL baselines
We investigate using four common RL algorithms: Deep Q-learning (dqn), Quantile Regression Deep Q-learning (qrdqn), Trust Region Policy Optimization (trpo), and Advantage Actor-Critic (a2c)—additional information on these algorithms is in Appendix D.2    RL Algorithm Descriptions. We use standard implementations of these methods available in the Stable-Baselines3 Python library (Raffin et al. 2021). To ground our analysis and discussion, note that dqn and qrdqn are off-policy methods which learn deterministic policies by directly solving for the optimal value function using the Bellman optimality condition in equation 2; exploration is induced only during training (e.g. epsilon-greedy). Whereas, trpo and a2c are on-policy methods which learn stochastic policies by direct optimization of the expected return in equation 1, using refinements of the policy gradient theorem (Sutton et al. 1999); exploration is inherent.
NWS and simple alternative baselines
To evaluate each of the RL policies, we compare them to the observed nws policy as well as several counterfactual baselines: most simplistically, randomly selecting $b$ days on which to issue alerts (random) and selecting the $b$ days with the highest QHI that summer (topk)—the latter is an oracle policy (not implementable in the real world) because to use it, we would have to know the whole summer’s daily QHI in advance. We also implement the general guidelines that NWS recommends in the absence of local criteria for issuing heat alerts (basic.nws): alert if the heat index is $\geq 100^{\circ}$ in northern states and $\geq 105^{\circ}$ in southern states (Hawkins, Brown, and Ferrell 2017). Lastly, we implement a policy of always issuing alerts on days above a per-county optimized QHI threshold until the budget runs out (aa.qhi).
Training data
The counties for which nws issued few heat alerts had high-variance estimates of heat alert effectiveness in the rewards model. Therefore, for the RL experiments, we selected a subset of 30 counties in which at least 75 heat alerts were issued during 2006-2016; spread across the five major climate regions, prioritizing those with higher variance in estimated alert effectiveness across days. See map in Figure 2 and Appendix D.1    Selection of 30 Counties for more details.
Evaluation metric
The main metric we use for policy comparison is the average cumulative reward per episode (“average return”). To estimate this metric fairly, we hold three years of data (2007, 2011, and 2015) out of RL training, referred to as the evaluation years [^6] Appendix D.4    Sensitivity Analysis Using 2014-2016 as the Evaluation Years shows the results of a sensitivity analysis where the models are trained on 2006-2013 and evaluated on 2014-2016.. For the purposes of RL hyperparameter tuning (details in Appendix D.3    Hyperparameter Tuning) and selection of the optimal QHI threshold for each county, we consider the regionally augmented weather trajectories and associated heat alert budgets from the evaluation years excluding the county of interest to be the validation data set. The final evaluation results (reported in tables and figures) are obtained by drawing 1,000 sets of coefficients from the Bayesian rewards model posterior and calculating the return under each policy with weather and budgets only from the county of interest during the evaluation years. These validation/tuning and evaluation procedures are described algorithmically in Appendix E    Summary of the Workflow Using BROACH. After calculating the average returns for the 30 counties under each competing policy, we compare each policy’s returns with the associated returns under the nws policy using a Wilcoxon-Mann-Whitney test. This nonparametric test allows us to compare the distribution of differences in returns, accounting for the fact that the counties are highly heterogeneous, so the differences between the policies are not normally distributed.
Sensitivity analysis
In the main analysis, we evaluate each county-specific implementation of dqn, qrdqn, trpo, and a2c by allowing each algorithm to deploy the type of policy that it optimizes during training: deterministic for dqn and qrdqn and stochastic for trpo and a2c. As a sensitivity analysis, we also deterministically evaluate the trpo and a2c policies, selecting whichever action has a higher probability under the policy function. The names of these models have the prefix “DET”. In another sensitivity analysis, we investigate whether the RL models can learn to anticipate subseasonal variation in the warm season by augmenting the RL state space with information about the future. Specifically, we test whether inclusion of the change in heat index over each of the next 10 days, as well as the 50th-100th (by every 10) percentiles of QHI for the remainder of the summer, improves RL performance. In real-time deployments of RL models, this kind of future information would not be known but could be sourced from weather forecasts or climate model projections. Of course, such forecasts / model projections are not perfect, but we start with perfect future information as a proof of concept. The names of models that include future information have the suffix “.F”.
5.3 Results and Discussion
Figures 3, S8, and S9 illustrate how the different policies look in practice, respectively in scenarios where using RL is beneficial, where the alert budget is small, and where using RL is not beneficial compared to the alternatives.
Figure 3: An example of observed and counterfactual heat alert policies for a single summer (2015 is the most recent year in our evaluation set), with estimates of the number of NOHR hospitalizations saved (compared to nws) per 10,000 Medicare enrollees under each policy. The dashed lines indicate the optimized QHI threshold of the policy in the same color. The multiple horizontal lines of pink and green dots indicate five different samples from trpo.qhi and a2c.qhi respectively. For these two policies, the number of NOHR hospitalizations saved is the average of all their evaluations over 2015.
Table 1 provides the main results of our RL experiments. (An extended version, including absolute number of hospitalizations, can be found in Table S6.)
Policy Median Difference WMW P-value
*topk 0.022 406 0.0002
random -0.015 177 0.88
basic.nws -0.286 30 1.0
aa.qhi 0.03 348 0.0025
dqn -0.123 43 0.99995
qrdqn -0.117 51 0.99991
trpo -0.065 97 0.99742
a2c -0.063 100 0.99689
dqn.qhi 0.03 370 0.00242
qrdqn.qhi 0.035 344 0.01121
trpo.qhi 0.038 338 0.0154
a2c.qhi 0.042 344 0.01121
dqn.f -0.417 40 0.99996
qrdqn.f -0.42 48 0.99993
trpo.f -0.062 103 0.99625
a2c.f -0.063 101 0.99669
dqn.qhi.f -1.895 1 1
qrdqn.qhi.f 0.014 238 0.45904
trpo.qhi.f 0.046 345 0.01062
a2c.qhi.f 0.036 343 0.01183
det.trpo.qhi -0.662 57 0.99985
det.a2c.qhi -0.886 90 0.99837
det.trpo.qhi.f 0.032 271 0.21723
det.a2c.qhi.f 0.02 287 0.13335
Table 1: Comparison between the average return of each counterfactual policy and that of the nws policy on the evaluation years, summarized across counties (e.g. “Median Diff.” is the median difference in average return). WMW is the Wilcoxon-Mann-Whitney statistic (higher is better); its associated p-value is also included. The first policy, marked by *, requires oracle knowledge of the future weather.
QHI restriction is necessary for the standard RL algorithms to perform well
Several counterfactual policies perform worse than nws (as indicated by negative median differences in returns): random, basic.nws, and the RL models without the QHI restriction. By contrast, the RL models with the QHI restriction perform significantly better than nws, as indicated by their significant positive median difference in returns. A rough estimate if a2c.qhi were implemented as-is across all counties in the U.S. (using the Medicare population from 2011, the midpoint of our study period) is that we could see a reduction of 222 NOHR hospitalizations per year (approximate 95% CI = (-491, 1131); details of this CI calculation are in Appendix F    Approximate Confidence Interval Calculation). Discussion of this absolute benefit to public health is in Appendix G    Discussion of Absolute Health Benefits.
The futility of including future information
Across both plain and QHI-restricted models, trpo was the only one that benefited from including future information in the state space. While a2c was minimally impacted, qrdqn and dqn were made noticeably worse. The latter may be due to the future information dramatically increasing the size of the state space/neural network parameters, which our hyperparameter tuning did not fully counteract. Note that for the trpo models, the benefit observed with the use of future information would likely be smaller in practice due to forecast / prediction error—the simulation of which is beyond the scope of this paper. Therefore, we focus on the RL models without future information in the remainder of this text.
The need for stochastic policies
Interpreting policies from the trpo and a2c models deterministically fails in the heat alerts setting under our rewards model. Additional discussion of this result is in Appendix H.2    Additional Discussion of Stochastic vs Deterministic Policies. This finding raises the question: Are stochastic policies palatable from a domain perspective? On the one hand, it is less immediately satisfying that some aspect of heat alerts issuance would be left to chance. On the other hand, in the context of human-in-the-loop decision making—which would likely be the case for public organizations such as NWS (Stuart et al. 2022)—an algorithm reporting probabilities is more informative than reporting only a binary action. In any case, if in the future a heat alerts RL model was running continuously online, it would likely need to utilize exploration (incorporating some randomness into its actions) to update itself over time.
Comparing policies that perform significantly better than NWS
We note that the performance and behavior of dqn.qhi is practically identical to aa.qhi—not just in its numerical summary, but also in its policy behavior. qrdqn.qhi performs a bit better than dqn.qhi, but not as well as the stochastic policies. Between those, a2c.qhi performs better than trpo.qhi, though their policies tend to display similar characteristics. Therefore, we focus on a2c.qhi as the best RL policy in the post-hoc analysis. The non-RL counterfactual policies that perform well are that of always issuing an alert on days above an optimized QHI threshold (aa.qhi) and of issuing alerts on the $b$ hottest days of the summer (topk). We note that the ordering of the Wilcoxon-Mann-Whitney statistics for topk, aa.qhi, and a2c.qhi is opposite the ordering of these policies’ median difference in returns (compared to nws). This is due to the longer tails of the a2c.qhi differences, as illustrated in Figure S11. For interpreting these results, recall that topk is an oracle policy. However, aa.qhi is an implementable alternative to a2c.qhi, so we must consider it more seriously.
5.4 Post-hoc Contrastive Explanation
To characterize differences between the best RL policies, the nws policy, and alternative baselines, we consider both stationary features included upstream in the analysis (e.g., climate region and median household income) as well as characteristics of both the nws and counterfactual alert policies, namely the distributions of the days of summer on which alerts are issued and the lengths of alert streaks (sequences of repeated alerts). We start with these alert characteristics, descriptive histograms of which are in Figure S10. Both a2c.qhi and aa.qhi tend to issue heat alerts earlier in the summer than nws. a2c.qhi also tends to issue shorter streaks of alerts (i.e., fewer consecutive days of alerts), which makes sense given its ability to learn sequential dependence encoded by the rewards model.
Figure S11 illustrates the varying performance of topk, aa.qhi, and a2c.qhi relative to the nws policy, across climate regions and counties. Note that large heterogeneity in the health effects of heat alert intervention across counties is consistent with previous findings (Wu et al. 2023). A striking visual pattern is the increasing vertical spread from left to right: topk performs the least heterogeneously across counties, followed by aa.qhi, and finally a2c.qhi. Similarly, the humid climate regions display more heterogeneity across counties than the others, across all the policies. This is partially due to those regions being larger and thus overrepresented among our 30 counties (see Table S4), but that does not fully explain the discrepancy.
To structure our investigation of why some counties see smaller (or even no) benefits from the application of the heat alerts RL, we fit CART on both the numeric difference in average return (compared to nws) and a polytomous indicator of which policy performed the best. To prevent CART overfitting to our 30 counties and to facilitate interpretation, we ensure that the minimum number of counties in each leaf (terminal) node is greater than or equal to five for the regression and three for the polytomous classification.
RL performs best in counties with larger heat alert-health signal and longer heat waves
A classification tree distinguishing a2c.qhi, aa.qhi, nws, and trpo.qhi is shown in Figure S12. The most predictive feature of the counterfactual policies’ performance relative to nws is the size of the alert budget. The fact that the other policies find it harder to improve on nws with more alerts may be because in such settings it is less critical to discriminate days when heat alerts will be most effective, given that our rewards model assumes more alerts never hurt in an absolute sense. The second split in the tree indicates that the RL performs best when there is greater variation in alert effectiveness across days, in other words, when there is more signal in the heat alert-health relationship that can be leveraged by the RL. The third split highlights how it is better to use RL in counties that tend to experience more prolonged heat waves, which cause aa.qhi to issue more alerts in a row, increasing the likelihood of alert fatigue. Running the same polytomous classification on a set of more conventional covariates (Figure S13), we see a nearly-identical tree except with median household income in place of alert effectiveness and humidity in place of prolonged heat waves. Both of these associations make sense from a domain science perspective.
Additionally, RL performs better than NWS earlier in the summer
Conducting regression on the difference in average returns between a2c.qhi and nws using CART (shown in Figure S14) generated complementary insights. The most predictive feature is the median day of summer on which the RL issued heat alerts: a2c.qhi performs better in counties for which it identified that it is optimal to issue alerts earlier in the summer. Whereas, in counties for which it is optimal to issue alerts later in the summer, there was less room for improvement by the RL because the nws already tends to issue alerts later in the summer.
6 Conclusion and Future Work Directions
This work lays the foundation of SDM for climate & health, by (1) formulating and building an SDM pipeline to optimize heat alert issuance, via integrating traditional statistical methods with cutting-edge RL techniques, and (2) offering insights into scenarios where RL improves vs. fails to improve on the observed policy or simpler alternatives.
Ultimately, we found that it was necessary to restrict alerts to days with higher QHI for standard RL algorithms (dqn, qrdqn, trpo and a2c) to be effective. Our post-hoc analysis enabled investigation of where RL performed better than the nws policy and the simpler alternative aa.qhi. Future work might explore the use of methods which ensure a new policy is never worse than the existing policy, such as safe policy learning or model predictive control. Alternatively, we could develop a preliminary predictive model to select counties that are likely to benefit from RL, using intuitions similar to those in our CART analysis.
We anticipate several common questions about our approach. The first is how to move beyond a fixed alert budget, important in a changing system (under climate change, the number of extreme heat events is expected to increase). From a practical standpoint, real monetary budgets from stakeholders could be used, for instance “how many times can we afford to open our cooling center(s) this summer?” Future work could also explore modeling alert fatigue as part of the RL environment, for instance by incorporating tools from behavioral science. RL algorithms that are intentionally robust to distribution shift also merit investigation in a climate & health setting.
A second question is whether offline (batch) RL methods, which use only observed data in place of a simulator, could be used to circumvent dependence on the specification of a rewards model. Several major challenges stand in the way of an offline approach, highlighted in our heat alerts example. First, there is a tension between trying to improve the observed policy and controlling distribution shift (Levine et al. 2020): for instance, we see that nws often issues alerts in streaks, making it hard for offline RL to learn and optimize the impact of an individual alert. Second, assessing the performance of offline RL using off-policy evaluation methods is made difficult by long episodes, large potential for distribution shift (especially in the absence of a modification such as restricting alerts to very hot days), and high degree of autocorrelation in the reward (Uehara, Shi, and Kallus 2022).
We close with a broader view of climate & health decision making. Ideally, an optimal heat alert system would account for different types of health impacts experienced by different demographic groups—future work could explore multi-objective RL. Similarly, it is important to recognize that the issuance of heat alerts is only a first step in reducing the public health impacts of extreme heat (Errett et al. 2023): there is much more work to be done analyzing and expanding local actions in response to heat alerts. If such on-the-ground interventions are able to increase the effectiveness of heat alerts, then this is likely to increase the ability of SDM methods such as RL to identify these effects and help us continue improving our strategies in the future.
Acknowledgements
This work was supported by an NSF Graduate Research Fellowship (EMC), NIH award K01ES032458 (RCN), NIH award R01-ES029950 (MT, GAW), NIH award 1R01MD016054-01A1 (FD), NIH award 5R01ES030616-04 (MT), NIH award 3RF1AG074372-01A1S1 (MT), NIH award P30ES000002 (MT), Alfred P. Sloan Foundation grant G-2020-13946 (FD), the Fernholz Foundation (MT), and Wellcome Trust grant 216033-Z-19-Z (GAW). We thank the National Studies on Air Pollution and Health research group for their support of this project. The computation in this paper was performed on the FASSE and FASRC Cannon clusters supported by the FAS Division of Science Research Computing at Harvard University. Lastly, we thank Kate Weinberger for facilitating access to the heat alerts data set, and the STAT 234 teaching staff at Harvard University for their feedback during early stages of this project.
Disclosures: GAW currently serves as a consultant for the Health Effects Institute (Boston, MA) and recently served as a consultant for Google, LLC (Mountain View, CA).
A    Heat Alerts MDP Terminology
Table S2 summarizes the heat alerts MDP components.
RL Term Notation Meaning in the Heat Alerts Application
episode $t\in\{0,\ldots,H-1\}$ the $H=152$ days of the warm season/summer in a specific year and county; $H$ is the horizon
state $s_{t}=(\xi_{t},x_{t})$ a vector of time-varying factors influencing the effect of heat on hospitalizations and the effectiveness of an alert; $\xi_{t}$ is the exogenous component (e.g., heat index) and $x_{t}$ is the endogenous component (e.g., history of heat alerts)
action $a_{t}$ whether a heat alert is issued (1) or not (0) at time $t$
reward $r_{t}=R(s_{t},a_{t})$ daily rate of heat-related hospitalizations at time $t$; a negative sign transform is applied so that fewer hospitalizations corresponds to a greater reward
agent the algorithmic decision-maker that implements a policy to determine when to issue heat alerts
policy $\pi(a_{t}|s_{t})$ the heat alert issuance rule; the NWS policy is the rule that was used to issue heat alerts in the U.S. from 2006-2016
cumulative reward / return $\sum_{h=t}^{H-1}r_{h}$ sum of remaining reward (transformed hospitalizations) at time $t$; unless otherwise specified, “cumulative” implies $t=0$
intervention / alert budget $\sum_{t=0}^{H-1}a_{t}\leq b$ the maximum number of heat alerts that can be issued in a given county-summer; in our experiments, the observed number of NWS alerts
Table S2: MDP/RL terminology in the context of the heat alerts environment.
B    Details of the Data Set
Evidence from the climate & health and environmental epidemiology literature guides our selection of relevant variables as well as feature engineering.
B.1    Heat Index and Heat Alerts
We follow Weinberger et al. (2021) and Wu et al. (2023) in using population-weighted maximum heat index to describe heat exposure in each county. They also combined NWS heat “warnings” and “advisories” into the single category of “alerts”. In some locations, where the forecast zones in which the NWS issues alerts are smaller than county boundaries, the alert history was taken from whichever forecast zone contained the largest proportion of the 2010 county population.
B.2    Medicare Hospitalizations
[…]
Figure S4 illustrates the rationale for excluding obviously heat-related hospitalization types from the pooled outcome, as described in Section 4.1.
Figure S4: Unobserved mediation of the effect of heat alerts on hospitalizations by heightened awareness.
B.3    Descriptive Statistics of Heat, Alerts, and Hospitalizations
Variable Min. Q1 Median Mean Q3 Max.
*NOHR hosps per 1,000 5.4 (3.1) 12.5 (12.7) 14.8 (15.2) 15 (15.3) 17.3 (17.8) 25.7 (34.5)
No. of NWS alerts 0 (0) 4 (0) 8 (3) 9.6 (4.8) 13 (7) 43 (57)
DMHI on alert days (°F) 70.5 (68.3) 100.3 (99.7) 103.5 (103.2) 102.5 (102.6) 106.2 (105.9) 121.1 (144)
Q-DMHI on alert days 18.9 (15.6) 91.6 (92.6) 95.8 (96.9) 93.4 (94.2) 98.2 (98.9) 100 (100)
No. of 90th pct. HI days 0 (0) 9 (8) 14 (14) 15.4 (15.4) 21 (21) 46 (61)
DOS of NWS alerts 16 (2) 68 (69) 84 (84) 83.4 (83.5) 99 (99) 142 (150)
DOS of 90th pct. HI days 15 (1) 72 (69) 88 (85) 86.9 (85.2) 102 (102) 145 (153)
Table S3: Summary statistics for our 30 counties of interest (and all the counties) of Medicare NOHR hospitalizations, NWS heat alerts, heat index, and day of summer for several features of interest. DMHI = daily maximum heat index, Q-DMHI = county quantile of DMHI (referred to simply as QHI in the main text), DOS = day of summer (out of 152). Variables given per county-summer are marked with *.
Table S3 provides descriptive statistics of the heat alerts dataset, both for the 30 counties in our RL analysis and all the counties with population $\geq$ 65,000 (used to create the BROACH environment). The two sets of statistics are similar aside from the number of heat alerts issued because the counties in the RL analysis were selected specifically for having issued higher numbers of alerts. Other observations are the long tail of the number of NWS alerts per county-summer, that some heat alerts were issued on days that were not very hot in an absolute or relative (quantile) sense, that there were some county-summers with few “hot” days (90th percentile by county [^7] Here, the 90th percentile was chosen to provide some intuition for the distribution of heat index and heat alerts. For context, Hondula et al. (2022) found that 95th percentile daily maximum heat index was the climatological indicator most strongly correlated with NWS heat alert frequency.) and others with many, and that the distribution of day-of-summer for NWS heat alerts and 90th percentile QHI by county are quite similar. A few other relevant statistics not shown in the table are that (across all the counties) there were a total of 40,387 heat alerts issued between 2006 and 2016, that 17.2% of NWS heat alerts were issued on days with heat index values that were not in their county’s top 90th percentile, and that 74.0% of days in the top 90th QHI by county did not have NWS heat alerts.
C    Details of the Bayesian Rewards Model
Recall our formulation for the expected per capita rate of NOHR hospitalizations at time $t$ in county-summer $(k,j)$ when taking action $a$:
$\rho_{t}^{(k,j)}(a)=\lambda_{k}(s_{t}^{(k,j)})(1-a\cdot\tau_{k}(s_{t}^{(k,j)}))$
The baseline and effectiveness functions
[…]
$\lambda(s_{t}^{(k,j)})$ $:=\exp\left(\beta_{k}^{\top}s_{t}^{(k,j)}\right),$ (5)
$\tau(s_{t}^{(k,j)})$ $:=\text{sigmoid}\left(\delta_{k}^{\top}s_{t}^{(k,j)}\right),$
[…]
Figure S5 illustrates the interaction of $\lambda_{k}$ and $\tau_{k}$ to capture the total effect of heat alerts on NOHR hospitalizations. For instance, we see that in early summer it is possible to have high heat alert effectiveness even when the heat index is not as high as it is later in the summer, possibly because people are less well-adapted to heat and/or more responsive to heat alerts early on.
Figure S5: Smoothed trajectories of observed quantile of heat index, modeled baseline NOHR hospitalization rate, and modeled alert effectiveness (assuming no past alerts) across days of summer; colored by climate region. Note that $\lambda$ and $\tau$ are normalized as they are in the RL reward function.
Data-driven prior
We propose a data-driven prior distribution that allows borrowing statistical strength from other locations while preserving the possibility of spatial heterogeneity. Let $w_{k}$ denote the vector of county-level features that are not included in the state vector but are predictive of the effect of heat and heat alerts on hospitalizations (see “County-level characteristics” in 4.1), and let $\gamma_{k}=[\beta_{k};\delta_{k}]=(\gamma_{k}^{\ell})_{\ell=1}^{L}$ denote the parameters corresponding to county $k$. Let N() denote a normal distribution. Then, the prior distribution of each $\gamma_{k}^{\ell}$ is set independently as $p_{\theta}(\gamma_{k}^{\ell}|\sigma_{\ell},w_{k})$ =
$\begin{cases}\mathrm{N}(\gamma_{k}^{\ell};f_{\theta}(w_{k})_{\ell},\sigma_{\ell}^{2})&\text{ no domain knowledge},\\ \mathrm{LogN}(\gamma_{k}^{\ell};\exp(f_{\theta}(w_{k})_{\ell}),\sigma_{\ell}^{2})&\text{ if }\gamma_{k}^{\ell}\in(0,\infty),\\ \mathrm{NegLogN}(\gamma_{k}^{\ell};-\exp(f_{\theta}(w_{k})_{\ell}),\sigma_{\ell}^{2})&\text{ if }\gamma_{k}^{\ell}\in(-\infty,0)\\ \end{cases}$ (6)
where $f_{\theta}$ is a feed-forward neural network (with weights $\theta$ and $L$ outputs) specifying the prior mean as a function of the spatial features, and $\sigma_{\ell}$ specifies the prior scale. We use a standard hyper-prior $p(\sigma_{\ell})=\mathrm{HalfCauchy(\sigma_{\ell})}$citeAgelman2006prior.
The domain knowledge constraints we utilize are: (i) past heat alerts cannot increase the baseline hospitalization rate—in other words, alert fatigue can be identified only in $\tau$—and (ii) higher QHI cannot decrease the effectiveness of heat alerts—note that this is conditional on day of summer.
[…]
Training
Traditional Bayesian inference methods using Markov Chain Monte Carlo are computationally prohibitive for our model due to the high dimensionality of parameters and the number of observations for each county. To address this issue, we use variational inference, which introduces an approximation of the true posterior distribution of the model parameters. We summarize the approach.
Let $(\gamma,\sigma)\sim q_{\psi}$ denote an approximation of the true posterior distribution where $\psi$ are the learnable parameters of the approximating distribution $q_{\psi}$. Next, let $o_{t}^{(k,j)}=(s_{t}^{(k,j)},a_{t}^{(k,j)},y_{t}^{(k,j)},n^{(k,j)})$ denote a data point in the training set and ${\mathcal{L}}(o_{t}^{(k,j)};\gamma_{k})$ be the Poisson log-likelihood corresponding to this data point (see equation 4). Variational inference attempts to maximize the likelihood of the observed data while penalizing the divergence between the variational posterior $q_{\psi}$ and the prior $p_{\theta}(\gamma,\sigma|w)={\prod_{k,l}}\,p_{\theta}(\gamma_{k}^{\ell}|\sigma_{\ell},w_{k})\,{\prod_{l}}\,p(\sigma_{\ell})$. The optimization objective is to maximize the evidence lower bound (ELBO) given by
$\textrm{ELBO}(\psi,\theta)$ $=\mathbb{E}_{q_{\psi}}\Big{[}{\textstyle{\sum_{k,j,t}}}{\mathcal{L}}(o_{t}^{(k,j)};\gamma_{k})\Big{]}$ (7)
$\>\>\>\>\>\>\>\>-D_{\mathrm{KL}}\left(q_{\psi}(\gamma,\sigma)\,\|\,p_{\theta}(\gamma,\sigma|w)\right),$
[…]
Results
The rewards model clearly converged in its loss across epochs (shown in Figure S6) and achieved higher predictive accuracy than standard machine learning algorithms (i.e., random forest and neural network). Predicting the rate of NOHR hospitalizations in the Medicare population in each county, this model has $R^{2}=0.102$—using the calculation $SSR/SST$ for comparability with earlier machine learning models (such as that we used when checking the signal in the mortality data). Predicting the absolute number of NOHR hospitalizations, it has $R^{2}=0.884$—much higher because the value is mostly driven by county population size. The model also displayed very good coverage when we ran it on synthetic data (1,000 samples from the posterior predictive) using known coefficients: average coverage across parameters for a 90% CI was 0.897.
Figure S6: Convergence of the ELBO in the Bayesian rewards model during training (80 epochs).
Figure S7 shows one posterior sample across the 761 counties on which the model was fit. Note that the spread across counties is larger than the spread within counties, so the IQR bars are no larger for multiple samples. Overall, we see that the coefficients on past alerts in today’s alert effectiveness ($\tau$) are both negative, empirical evidence of alert fatigue. In the baseline hospitalization rate ($\lambda$), we see a strong protective effect by weekend.
Figure S7: A sample of posterior coefficients from the rewards model (in $\tau$ and $\lambda$ respectively), displaying IQR and median. Bias and coefficients for the spline on day of summer are not shown, for ease of interpretability. (+)/(-) indicates a coefficient constrained to be positive/negative.
D    Details of the RL Analysis
D.1    Selection of 30 Counties
County Fips (State name abbreviation) Region No. Alerts (Eval Years) Population Density (/mi2) Median HH Income (USD) Democratic Voters (%) Broadband Usage (%) Average PM2.5 ($\mu g/m^{3}$)
5045 (AR) MxHd 61 179.00 50314.00 34.00 37.50 9.08
20161 (KS) MxHd 47 120.00 43962.00 43.50 52.20 6.35
21059 (KY) MxHd 24 212.00 46555.00 38.70 53.30 10.22
29019 (MO) MxHd 50 242.00 48627.00 51.60 46.60 7.85
37085 (NC) MxHd 31 200.00 44625.00 39.30 46.70 9.65
40017 (OK) MxHd 63 133.00 63629.00 22.80 70.10 7.53
47113 (TN) MxHd 58 176.00 41617.00 44.30 79.40 9.29
42017 (PA) MxHd 25 1036.00 76555.00 51.00 85.10 9.14
41053 (OR) Mrn 31 102.00 52808.00 45.60 79.20 3.24
41067 (OR) Mrn 31 745.00 64180.00 57.70 82.70 4.28
53015 (WA) Mrn 23 90.00 47596.00 49.40 53.60 3.59
13031 (GA) HtHd 43 106.00 35840.00 38.90 57.50 9.01
22109 (LA) HtHd 34 91.00 49960.00 27.90 55.50 8.26
28049 (MS) HtHd 57 283.00 37626.00 69.90 36.50 8.50
45015 (SC) HtHd 48 168.00 52427.00 41.20 66.20 8.72
48157 (TX) HtHd 42 707.00 85297.00 47.80 83.70 8.64
48367 (TX) HtHd 41 131.00 64515.00 18.30 52.20 7.84
22063 (LA) HtHd 42 201.00 56811.00 13.70 47.60 8.87
4015 (AZ) HtDr 29 15.00 39200.00 28.70 63.60 4.59
6025 (CA) HtDr 40 42.00 41807.00 64.00 64.40 7.22
32003 (NV) HtDr 29 251.00 52873.00 55.90 78.10 5.21
4013 (AZ) HtDr 72 423.00 53596.00 44.00 77.40 6.38
6071 (CA) HtDr 29 103.00 54090.00 51.60 77.10 6.38
17115 (IL) Cold 22 190.00 46559.00 45.70 63.60 8.87
17167 (IL) Cold 25 228.00 55449.00 46.00 66.50 8.75
19153 (IA) Cold 24 764.00 59018.00 54.90 57.90 8.25
19155 (IA) Cold 27 98.00 51304.00 44.30 42.00 7.89
34021 (NJ) Cold 29 1639.00 73480.00 67.30 81.60 9.31
29021 (MO) Cold 44 219.00 44363.00 43.70 48.80 7.97
31153 (NE) Cold 27 681.00 69965.00 37.60 72.60 7.92
Table S4: Descriptive statistics of the 30 counties used in policy optimization. Climate region abbreviations = {MxHD: Mixed-Humid, Mrn: Marine, HtHd: Hot-Humid, HtDr: Hot-Dry, Cold: Cold}. HH = household. PM2.5 is an annual average.
Out of the 170 counties with 75 or more heat alerts issued between 2066 and 2016, we selected 30 using the following criteria: for each climate region [^8] Excluding “Mixed Dry” and “Very Cold” because there were very few counties in these regions with sufficient population size, select the county with the highest variance in estimated heat alert effectiveness (indicating higher signal as identified by our rewards model). To get up to 30, repeat this selection, but first prioritize selecting a county from a U.S. state not already in the set, and if we must repeat a state, do not select a county that is directly adjacent to one already in the set. If there are no more eligible counties in a region, select from a larger region. Characteristics of these 30 counties are provided in Table S4.
D.2    RL Algorithm Descriptions
[…]
D.3    Hyperparameter Tuning
[…]
D.4    Sensitivity Analysis Using 2014-2016 as the Evaluation Years
For real-world deployment of a heat alert-RL system, an ideal setup would be to train on years $\{1,...,x\}$ and evaluate on year $x+1$. However, in this analysis we held out $\{2007,2011,2015\}$ as the evaluation set to provide a better representation of historical performance. Note that while days within each episode are consecutive, the 7-month gap between episodes lessens concern about correlated episodes.
Evaluation Period QHI Median Median Diff. (RL-NWS) WMW p-value
Main Analysis:
{2007, 2011, 2015} 0.042 0.011
2007 0.51 0.006 0.013
2011 0.555 0.016 0.027
2015 0.519 0.01 0.010
Sensitivity Analysis:
{2014, 2015, 2016} 0.007 0.121
2014 0.46 0.002 0.012
2015 0.519 0.009 0.022
2016 0.6 -0.005 0.777
Table S5: Comparison of results using 2014-2016 as the evaluation years (sensitivity analysis) as opposed to 2007,2011,2015 (main analysis).
To provide intuition for how sequential training-testing would look, we re-ran the best RL model (A2C.QHI) using 2006-2013 to train and 2014-2016 to evaluate. To reduce the computational burden, we used the optimized hyperparameters from the main analysis for each county. Table S5 shows the aggregated results as well as results stratified by year. A few notable results are that (i) the ordering of the original results ($2007<2015<2011$) is the same as the ordering of median QHI, (ii) the 2015 results are similar in each evaluation scheme, and (iii) the 2016 result is the worst. However, this last result is overly bleak because in an online deployment we would not be evaluating 3 years ahead—the model would be updated at least once a year.
E    Summary of the Workflow Using BROACH
This section is meant to illustrate a toy example of our workflow. Imagine seven counties located in two climate regions, as shown in the figure below.
Let the training years TY = $\{2006,2008,2009,2010,2012,2013,2014,2016\}$ and evaluation years EY = $\{2007,2011,2015\}$.
First, train the rewards model on counties 1-7 and years TY $\cup$ EY to obtain posterior distributions for $r_{1}(s,a),\>...,r_{7}(s,a)$. Then, for each county, follow the steps below. Consider county 1 for illustration. Our procedure is as follows:
1:if Training RL (with a QHI restriction) then
[…]
3:         for i in 1 to $N_{T}$ episodes do
4:              Sample $r_{1,i}(s,a)$ from the posterior for $r_{1}(s,a)$
5:              Sample $\xi_{i}\sim TY\times\{\xi_{1},\xi_{2},\xi_{3},\xi_{4}\}$
6:              Update $\pi_{1,h}(a|s)$ s.t. a = 0 if QHI $<$ h, using $r_{1,i}$ and $\xi_{i}$
7:         end for
8:         Select $\pi^{*}_{1,h}$ (across i) with the largest average return using $\xi_{i}\sim EY\times\{\xi_{2},\xi_{3},\xi_{4}\}$
9:     end for
10:else if Tuning RL or aa.qhi then
[…]
12:         for i in 1 to $N_{E}$ episodes do
13:              Sample $r_{1,i}(s,a)$ from the posterior for $r_{1}(s,a)$
14:              Sample $\xi_{i}\sim EY\times\{\xi_{2},\xi_{3},\xi_{4}\}$
15:              Calculate cumulative reward using $\pi^{*}_{1,h}$ (or aa.qhi with h), $r_{1,i}$ and $\xi_{i}$
16:         end for
17:     end for
18:     Select h with the largest return
19:else if Final evaluation for any policy (but using notation for optimized RL) then
20:     for i in 1 to $N_{E}$ episodes do
21:         Sample $r_{1,i}(s,a)$ from the posterior for $r_{1}(s,a)$
22:         Sample $\xi_{i}\sim EY\times\{\xi_{1}\}$
23:         Calculate cumulative reward using $\pi^{*}_{1,h^{*}}$, $r_{1,i}$ and $\xi_{i}$
24:     end for
25:end if
F    Approximate Confidence Interval Calculation
To generate an approximate confidence interval for the absolute number of NOHR hospitalizations saved by implementing our best RL model, we use the estimates from the final evaluation for a2c.qhi and nws. However, rather than first calculating average returns per county (across the 1,000 evaluation episodes) and then taking the median across the 30 counties, we first take the median return (across the 30 counties) from each of the 1,000 evaluation episodes and then calculate the 2.5% and 97.5% quantiles of the resulting distribution of medians.
G    Discussion of Absolute Health Benefits
Do we estimate substantive health benefits from implementing the improved heat alerts policies detailed in the main text? For general context, the median rate of NOHR hospitalizations per 10,000 Medicare enrollees per summer during our study period was 148 (Table S3), and we estimated that deploying our best RL model relative to issuing no heat alerts would save 2 of these 148, or 1.4%, in the 30 counties in our RL analysis. Noting that the overall relative risk of a Medicare enrollee experiencing one of the NOHR hospitalization types on a heat wave day is about 1.10 (Bobb et al. 2014), this is substantial evidence in support of continuing to issue heat alerts to protect public health.
Compared to the observed nws policy, a2c.qhi applied to the 30 counties in our analysis would save approximately 0.042 NOHR hospitalizations per 10,000 Medicare enrollees per summer, a reduction of only 0.03%. Multiplying this by 49 million, the size of the Medicare population in 2011 (the midpoint of our study period), yields about 222 NOHR hospitalizations saved per summer. However, this modest value should be further contextualized by several points.
Firstly, heat alerts by themselves (and in our case, the re-arrangement of the same number of heat alerts) are a very cost-effective intervention, so even small reductions in health harms are promising.
Secondly, if we take into account the large heterogeneity in the benefit of RL across counties with numbers from the CART regression (Figure S14), assuming that we figure out how to implement a safe policy such that counties which would not benefit are unaffected, this number increases to 262 ((49m/10k)*(0.074*0.17 + 0.053*0.33 + 0.18*0.13)) NOHR hospitalizations saved per summer. We obtain a nearly identical result from running a CART analysis using only variables which are available for all 761 counties that were included in the rewards model, and applying the results of that model across the whole set of counties. These are rough approximations given that they assume the proportion of counties in which RL alert policies were found to improve health outcomes in our sample is generalizable to the rest of the US. It is plausible that our sample of 30 counties is unrepresentative of the U.S. as they tend to have higher variability in alert effectiveness (increasing potential improvements from RL) and also higher alert budgets.
[…]
Fourthly, the numbers reported in this study should not be interpreted as the overall benefit to public health from improving heat alerts issuance strategies. The health data used in our study only reflect inpatient hospitalizations from NOHR causes billed to Medicare Part A, or hospital fee-for-service. Only about 60% of Medicare enrollees are in the fee-for-service population [^9] https://www.cms.gov/Research-Statistics-Data-and-Systems/Statistics-Trends-and-Reports/Beneficiary-Snapshot/Downloads/Bene˙Snaphot.pdf, so hospitalizations of the other 40% do not appear in our dataset.
H    Extended Results from the RL Analysis
H.1    Policy Examples
Figure S8: Example of a setting with relatively few heat alerts.
Figure S9: Example of a setting where implementing RL would not be beneficial relative to nws or aa.qhi.
H.2    Additional Discussion of Stochastic vs Deterministic Policies
Interpreting policies from the trpo and a2c models deterministically fails because many of the policy functions’ estimated probabilities for issuing an alert are below 0.5, so always taking the action with the larger probability dramatically reduces the number of alerts issued—in some counties to zero. Note that policies which do not use all the alerts in their budget are penalized by the specification of our rewards model, resulting in det.trpo.qhi and det.a2c.qhi no longer outperforming nws with statistical significance across the 30 counties. Inclusion of future information mitigates this issue, but the results are not nearly as good as their stochastic counterparts.
Our big picture intuition for this result is that we are in a heavily data-constrained setting, which frustrates RL’s ability to completely identify alert fatigue. The stochastic models address this by sampling alert days, inducing less alert streaks. Thought a bit unsatisfying, we believe this is an acceptable compromise given that, as mentioned in the main text, an online heat alert-RL system would likely utilize stochastic exploration anyway, to update itself over time.
NOHR Hosps Saved vs NWS (vs Zero) Per Summer
Policy Median Diff. WMW P-value Median / 10,000 Medicare Approximate Total
*topk 0.022 406 0.0002 0.023 (1.96) 113 (9,603)
random -0.015 177 0.88 -0.013 (1.98) -64 (9,697)
basic.nws -0.286 30 1.0 -0.283 (1.45) -1,388 (7,110)
aa.qhi 0.03 348 0.0025 0.029 (1.99) 143 (9,747)
dqn -0.123 43 0.99995 -0.102 (1.83) -502 (8,989)
qrdqn -0.117 51 0.99991 -0.098 (1.85) -480 (9,046)
trpo -0.065 97 0.99742 -0.063 (1.9) -309 (9,321)
a2c -0.063 100 0.99689 -0.062 (1.9) -301 (9,317)
dqn.qhi 0.03 370 0.00242 0.029 (1.99) 142 (9,746)
qrdqn.qhi 0.035 344 0.01121 0.031 (1.99) 153 (9,767)
trpo.qhi 0.038 338 0.0154 0.038 (2.02) 185 (9,897)
a2c.qhi 0.042 344 0.01121 0.045 (2.02) 222 (9,902)
dqn.f -0.417 40 0.99996 -0.401 (0.49) -1,966 (2,408)
qrdqn.f -0.42 48 0.99993 -0.42 (0.47) -2,059 (2,322)
trpo.f -0.062 103 0.99625 -0.061 (1.9) -297 (9,322)
a2c.f -0.063 101 0.99669 -0.062 (1.9) -302 (9,304)
dqn.qhi.f -1.895 1 1 -1.901 (0) -9,317 (0)
qrdqn.qhi.f 0.014 238 0.45904 0.012 (1.61) 61 (7,876)
trpo.qhi.f 0.046 345 0.01062 0.048 (2.02) 236 (9,886)
a2c.qhi.f 0.036 343 0.01183 0.038 (2.02) 187 (9,915)
det.trpo.qhi -0.662 57 0.99985 -0.579 (0) -2,835 (2)
det.a2c.qhi -0.886 90 0.99837 -0.794 (0.16) -3,888 (807)
det.trpo.qhi.f 0.032 271 0.21723 0.034 (1.83) 168 (8,948)
det.a2c.qhi.f 0.02 287 0.13335 0.021 (2.01) 104 (9,836)
[…]
H.3    Additional Figures from the Post-hoc Contrastive Analysis
Figure S10 provides histograms of different policies’ temporal characteristics, figure S11 compares different policies’ performance (compared to NWS) across climate regions and counties, and figures S12, S13, and S14 illustrate the CART analysis.
Figure S10: Distributions of heat alert characteristics under different policies, over the evaluation years.
Figure S11: Different policies’ average return relative to nws across different climate regions, over the evaluation years. Outlier points are labeled with their state abbreviation. Note that topk is an oracle policy.
Figure S12: Polytomous classification tree comparing a2c.qhi, aa.qhi, nws, and trpo.qhi across the 30 counties. Within each node, the policy at the top is that associated with the highest probability of performing best. The four probabilities beneath correspond to the four policies in alphabetical order, as written above. (Note that the two RL policies tend to be more similar to one another, as do the aa.qhi and nws policies, so the distribution of probabilities is important to consider in addition to the overall classification of each node.) The percentage at the bottom is the fraction of counties represented by that node. “No. Alerts” is the average number of heat alerts per warm season during the evaluation years, “Alert Effectiveness” is the $\tau$ estimated by our rewards model, and SL is the streak length.
Figure S13: Polytomous classification tree comparing a2c.qhi, aa.qhi, nws, and trpo.qhi across the 30 counties—using more conventional features than Figure S12. Within each node, the policy at the top is that associated with the highest probability of performing best. The four probabilities beneath correspond to the four policies in alphabetical order, as written above. (Note that the two RL policies tend to be more similar to one another, as do the aa.qhi and nws policies, so the distribution of probabilities is important to consider in addition to the overall classification of each node.) The percentage at the bottom is the fraction of counties represented by that node. “No. Alerts” is the average number of heat alerts per warm season during the evaluation years.
Figure S14: Regression tree on the differences in average return between a2c.qhi and nws. Within each node, the number at the top is the prediction (average) difference, with positive indicating that the RL is better and negative indicating that the nws is better. The percentage at the bottom is the fraction of counties represented by that node. DOS is the day of summer, SL is the streak length, and “No. Alerts” is the average number of heat alerts per warm season during the evaluation years.
[…]